# Interpreting the Effects of Quantization on LLMs

### v1

@@ Abstract
Quantization offers a practical solution to deploy LLMs in resource-constraint environments. However, its impact on internal representations remains understudied, raising questions about the reliability of quantized models. In this study, we employ a range of interpretability techniques to investigate how quantization affects model and neuron behavior. We analyze multiple LLMs under 4-bit and 8-bit quantization. Our findings reveal that the impact of quantization on model calibration is generally minor. Analysis of neuron activations indicates that the number of dead neurons, i.e., those with activation values close to 0 across the dataset, remains consistent regardless of quantization. In terms of neuron contribution to predictions, we observe that smaller full precision models exhibit fewer salient neurons, whereas larger models tend to have more, with the exception of Llama-2-7B. The effect of quantization on neuron redundancy varies across models. Overall, our findings suggest that effect of quantization may vary by model and tasks, however, we did not observe any drastic change which may discourage the use of quantization as a reliable model compression technique.
@@ 1 Introduction
The last decade has seen a tremendous amount of work done in language modeling, specifically in large language models (LLMs) Devlin et al. (2019); Liu et al. (2023a); Touvron et al. (2023). There is a common trend to increase the number of parameters in LLMs to improve performance. However, this approach exacerbates the challenge of resource requirements, including computational and energy costs Patterson et al. (2021). Quantization is a model compression technique that is widely used because of its effectiveness and simplicity Bondarenko et al. (2024); Dettmers et al. (2022); Wu et al. (2023). Quantization reduces the model size by using lower precision weights and/or activations, which can improve its inference speed while using less storage space. The effect of quantization is generally measured by comparing a model’s performance on downstream NLP tasks Li et al. (2024); Kurtić et al. (2024).
While performance on downstream tasks is crucial to understand the end-to-end impact, the evaluation is limited to a set of downstream tasks used for evaluation. In other words, it does not provide complete insights into the effect of quantization on the knowledge learned by models. In this work, we argue that the interpretation serves as an additional metric and evidence to analyze the effect of quantization on the model. For instance, it may reveal which types of knowledge or relationships are preserved or degraded by quantization, giving a deeper understanding of whether essential patterns remain intact. This is especially important for safety-critical applications such as finance, law, and healthcare Hassan et al. (2024) where reliability of a model is necessary.
@@
    Does quantization influence the contribution of neurons to model predictions?
@@
    How does quantization affect the number of “dead neurons”?
@@
    Does quantization affect the redundancy of neurons? In other words, does it result in more neurons learning identical information?
We analyze multiple open-source models, under two quantization settings: 4-bit Dettmers et al. (2023) and 8-bit Dettmers et al. (2022) and compare them with the full-precision float-16 weight model.
@@ 2.1 Confidence Analysis
Confidence analysis aims to find the average confidence of a model in its predictions Abdar et al. (2021). We calculate the average confidence of the model using the following equation:
@@ 2.2 Calibration Analysis
Calibration can be defined as the degree to which a model’s predicted probabilities reflect the actual frequencies of those outcomes Nixon et al. (2020). Despite high accuracy, deep neural networks often suffer from miscalibration Guo et al. (2017).
We use the Adaptive Calibration Error (ACE) metric Nixon et al. (2020), which adjusts its assessment based on the actual distribution of confidence values, enabling a more flexible and precise evaluation of calibration. ACE is calculated as follows:
@@ 2.3 Neuron’s Attribution
A neuron’s attribution refers to its role and significance in a model’s predictions, as determined by attribution methods such as integrated gradient (IG) Sundararajan et al. (2017). To evaluate the impact of quantization on neuron attributions, we analyze the number of salient neurons that contribute significantly to the model’s predictions. This analysis shows quantization effects on the model’s ability to identify and rely on the important features.
@@
Most attributed token-based: In this technique, we only consider the most attributed token’s (i.e., input token with max attribution score) representation and select neurons that have a normalized attribution score >0.8>0.8. This identifies neurons that are most important in determining the model’s predictions for the specific context of the selected token. Given as:
@@
$n_{j}^{\text{salient}}=\{n_{j}\mid\frac{n_{j}}{\max(n_{j})}>0.8\},\forall j\in\text{Layer}$
@@
Input sequence-based: To identify neurons that are salient in the context of the input sequence, we calculate the total attribution over the entire input sequence by summing the attributions across all input tokens. We select the neurons that have an attribution score >0.8>0.8 after normalization. This approach ensures that the selected neurons reflect their contributions to the overall meaning of the input, rather than being limited to the most attributed token only. Given as:
@@
$n_{j}^{\text{salient}}=\{n_{j}\mid\frac{s_{j}}{\max(s_{j})}>0.8\},\forall j\in{Layer}$
@@
$n_{j}^{\text{salient}}=\{n_{j}\mid\frac{m_{j}}{\max(m_{j})}>0.8\},\forall j\in{Layer}$
@@ 2.4 Neuron’s Activations
Since quantization reduces weight precision, it may increase the number of insignificant neurons. To identify them, we follow Voita et al. (2023), defining dead neurons as those whose activations remain consistently near zero across the dataset.
@@ 2.4.1 Dead/Insignificant Neurons
Voita et al. (2023) observed that the number of dead neurons increases with the growth of a model’s size. Their analysis of the OPT language model family, which uses the ReLU activation function, shows that over 70% of neurons in some layers are dead. We hypothesize that quantization, by reducing the precision of weights, may contribute to an increase in the number of dead neurons in the network.
Apart from ReLU, other activation functions such as GELU Hendrycks and Gimpel (2016) and SiLU Elfwing et al. (2017) may not produce activation values that are exactly zero. To generalize the concept of dead neurons for these activation functions, we define a threshold of 0.1-0.1 to $0.1$, categorizing neurons as dead if their activation values consistently remain within this range across the dataset. For different activation functions, we define dead neurons as follows:
@@ 2.5 Correlation Analysis
We hypothesize that a low-precision quantization may cause more neurons to represent identical information, i.e., as precision is reduced, high precision neuron values may map to the same low precision value. Similar to Dalvi et al. (2020), we calculate the Pearson correlation of neurons at a layer to identify neurons representing similar information. The Pearson correlation is given by:
@@ 3.1 Datasets
We consider five datasets in this study: BoolQ Clark et al. (2019), the Jigsaw Toxicity dataset cjadams et al. (2017), Physical Interaction: Question Answering (PIQA) Bisk et al. (2020), Hellaswag Zellers et al. (2019) and IMDB sentiment classification Maas et al. (2011).
We select a random subset of the each dataset for our experiment. More specifically, we used 10k samples from a combination of train and validation sets of BoolQ, 9k samples from the Toxicity train set, 1,838 validation samples from PIQA, 5,000 validation samples from Hellaswag, and the IMDB training set. Instruction-tuned samples for datasets is available in Appendix A. We used NeuroX Dalvi et al. (2023) for activations extraction.
@@ 3.2 Models
The primary models analyzed in our study are Phi-2 Javaheripi and Bubeck (2023), Llama-2 7B Touvron et al. (2023), Qwen 2.5 3B and 7B Qwen et al. (2025), and Mistral-7B Jiang et al. (2023).
@@
Since our subject models employ GELU (Phi-2) and SiLU (Llama, Qwen, Mistral) as activation functions, which do not produce exact zero activations, we include the OPT-6.7B model from the OPT family Zhang et al. (2022) to assess the behavior of ReLU activations for comparison. This model utilizes a decoder-only architecture similar to other subject models.
@@ 3.3 Quantization Configurations
To perform comparative analysis across models under different quantization settings, we employed two widely-used quantization techniques: 4-bit Dettmers et al. (2023) and 8-bit Dettmers et al. (2022). Models are quantized using bitsandbytes config through Huggingface transformers. Table 1 shows the hyperparameters used for quantization.
@@ 3.4 Attribution Technique
We use Integrated Gradients Sundararajan et al. (2017) using Captum Kokhlikyan et al. (2020) to find salient neurons in a network.
@@ 4.2.2 Calibration Analysis
The seemingly random fluctuations in ACE scores, particularly for certain model, dataset or weight precision combinations, could stem from several underlying factors. Differences in model pretraining objectives strategies may contribute to how calibration responds to quantization. Although quantization may introduce fluctuations in ACE, the difference is not substantial to undermine it’s reliability, even often yielding improved calibration relative to full-precision variant.
@@ 4.3 Effect on the Contribution of Neurons to Model Predictions
Overall, the number of salient neurons serves as a proxy for how distributed or localized the decision-making process is within the network Dalvi et al. (2020); Durrani et al. (2024). Full precision models tend to use fewer neurons, when they are smaller. In contrast, in larger models, full precision can enable richer and more distributed neuron contribution.
4.4 Effect on the number of “dead neurons”
@@
As shown in Table 3, quantization causes only a minor change in the count of dead neurons. The trend across quantization seems to be consistent, as the number of dead neurons remains almost similar between quantized and full-precision models. The pattern of higher neurons in initial layer in Phi-2 and OPT-6.7B likely reflects the role of initial layers in learning sparse, low-level features, while later layers capture higher-level contextual features Dalvi et al. (2022); Voita et al. (2023). We hypothesize that the consistently low count of dead neurons among Llama, Qwen and Mistral is due to the use of the SiLU activation function.
@@ 4.5 Effect on the Redundancy of Neurons
As identified in the works of Dalvi et al. (2020) language models can maintain 97% of accuracy while using only 10% of the neurons. This finding is valuable for model pruning. We investigate whether quantization leads to higher redundancy. Due to the substantial computational requirements, our analysis was limited to the Phi-2 and Llama-2-7B models, using activations from BoolQ.
@@ 4.5.1 Correlation Analysis
Figure 4 shows the number of neuron pairs per correlation bin for 4-bit, 8-bit, and full-precision variants of Phi-2 and Llama-2-7B. The X-axis represents correlation bins (0.3-0.4 to 0.9-1.0), and the Y-axis shows the total count of neuron pairs across all layers. This binning highlights redundancy across quantization levels. For clarity, bins below 0.3 were omitted, as they showed similar counts of uncorrelated neurons across all variants.
@@ 5 Related Work
Quantization Techniques. Quantization Gray and Neuhoff (1998) is used to reduce the memory requirement by reducing the size of weight and/or activation and increasing the inference time of a model Jacob et al. (2017); Gholami et al. (2021).
Quantization-aware training (QAT) is costly and uses re-training of a model on a dataset to maintain accuracy Liu et al. (2023b); Du et al. (2024); Dettmers et al. (2023); Kim et al. (2023).
Post-training quantization quantizes models without any additional finetuning of the model with a limited dataset, but also suffers from performance issues Banner et al. (2019); Cai et al. (2020). In case of LLM’s Post Training Quantization can be of 3 types: i) Weight-Only Quantization Park et al. (2024); Frantar et al. (2023); Chee et al. (2024); Lin et al. (2024), ii) Weight-Activation Quantization Yao et al. (2022); Yuan et al. (2023); Guo et al. (2023); Wei et al. (2023), and iii) KV Cache Quantization Hooper et al. (2024); Yue et al. (2024).
Quantization Analysis and Interpretation. Xia et al. (2021) explores confidence and calibration relation between quantized and full-precision model by using symmetric quantization. Proskurina et al. (2024) shows quantization improves calibration in LLMs using GPTQ. Some literature exlores interpretation withing quantized model for vision model Norrenbrock et al. (2024); Arazo et al. (2024); Maleki et al. (2024); Rezabeyk et al. (2024); Amine KERKOURI et al. (2024).

### v2

@@ Abstract
Quantization offers a practical solution to deploy LLMs in resource-constraint environments. However, its impact on internal representations remains understudied, raising questions about the reliability of quantized models. In this study, we employ a range of interpretability techniques to investigate how quantization affects model and neuron behavior. We analyze multiple LLMs under 4-bit and 8-bit quantization. Our findings reveal that the impact of quantization on model calibration is generally minor. Analysis of neuron activations indicates that the number of dead neurons, i.e., those with activation values close to 0 across the dataset, remains consistent regardless of quantization. In terms of neuron contribution to predictions, we observe that smaller full precision models exhibit fewer salient neurons, whereas larger models tend to have more, with the exception of Llama-2-7B. The effect of quantization on neuron redundancy varies across models. Overall, our findings suggest that effect of quantization may vary by model and tasks, however, we did not observe any drastic change which may discourage the use of quantization as a reliable model compression technique. We make our code publicly available. [^1] https://github.com/MSingh-CSE/LLM-Quantization-Effects
@@ 1 Introduction
The last decade has seen a tremendous amount of work done in language modeling, specifically in large language models (LLMs) devlin2019bertpretrainingdeepbidirectional; Liu_2023; touvron2023llama2openfoundation. There is a common trend to increase the number of parameters in LLMs to improve performance. However, this approach exacerbates the challenge of resource requirements, including computational and energy costs patterson2021carbonemissionslargeneural. Quantization is a model compression technique that is widely used because of its effectiveness and simplicity bondarenko2024lowrankquantizationawaretrainingllms; dettmers2022llmint88bitmatrixmultiplication; wu2023understandingint4quantizationtransformer. Quantization reduces the model size by using lower precision weights and/or activations, which can improve its inference speed while using less storage space. The effect of quantization is generally measured by comparing a model’s performance on downstream NLP tasks li2024evaluatingquantizedlargelanguage; neuralmagicquantization.
While performance on downstream tasks is crucial to understand the end-to-end impact, the evaluation is limited to a set of downstream tasks used for evaluation. In other words, it does not provide complete insights into the effect of quantization on the knowledge learned by models. In this work, we argue that the interpretation serves as an additional metric and evidence to analyze the effect of quantization on the model. For instance, it may reveal which types of knowledge or relationships are preserved or degraded by quantization, giving a deeper understanding of whether essential patterns remain intact. This is especially important for safety-critical applications such as finance, law, and healthcare hassan2024activelearningrobustrepresentative where reliability of a model is necessary.
@@
    Does quantization influence the contribution of neurons to model predictions? A major change in contributing neurons may reflect a change in model’s decision making process.
@@
    How does quantization affect the number of “dead neurons”? A representation with a large number of dead neurons can cause the model to depend on only a handful of neurons.
@@
    Does quantization affect the redundancy of neurons? In other words, does it result in more neurons learning identical information which may also reflect that some learned knowledge is lost during quantization.
Understanding how quantization affects a model’s confidence and calibration is essential, as distortions in output probabilities can lead to unreliable uncertainty estimates in critical applications. Examining whether quantization alters the contribution of neurons to predictions provides insight into how compression reshapes neurons sensitivity to input features. Analyzing changes in the number of dead neurons helps assess whether quantization reduces network utilization or representational capacity, which can undermine robustness. Finally, exploring the redundancy of neurons i.e. whether quantization leads to more neurons learning similar information, offers valuable understanding of how model compression influences feature diversity and efficiency. Collectively, these objectives aim to reveal how quantization not only affects performance metrics but also the deeper structural and behavior of LLMs.
We analyze multiple open-source models, under two quantization settings: 4-bit dettmers2023qloraefficientfinetuningquantized and 8-bit dettmers2022llmint88bitmatrixmultiplication and compare them with the full-precision float-16 weight model.
@@ 2.1 Confidence Analysis
Confidence analysis aims to find the average confidence of a model in its predictions 10.1016/j.inffus.2021.05.008. We calculate the average confidence of the model using the following equation:
@@ 2.2 Calibration Analysis
Calibration can be defined as the degree to which a model’s predicted probabilities reflect the actual frequencies of those outcomes nixon2020measuringcalibrationdeeplearning. Despite high accuracy, deep neural networks often suffer from miscalibration guo2017calibrationmodernneuralnetworks.
We use the Adaptive Calibration Error (ACE) metric nixon2020measuringcalibrationdeeplearning, which adjusts its assessment based on the actual distribution of confidence values, enabling a more flexible and precise evaluation of calibration. ACE is calculated as follows:
@@ 2.3 Neuron’s Attribution
A neuron’s attribution refers to its role and significance in a model’s predictions, as determined by attribution methods such as integrated gradient (IG) sundararajan2017axiomatic. To evaluate the impact of quantization on neuron attributions, we analyze the number of salient neurons that contribute significantly to the model’s predictions. This analysis shows quantization effects on the model’s ability to identify and rely on the important features.
@@
Most attributed token-based: In this technique, we only consider the most attributed token’s (i.e., input token with max attribution score) representation and select neurons that have a normalized attribution score >0.7>0.7. This identifies neurons that are most important in determining the model’s predictions for the specific context of the selected token. Given as:
@@
$n_{j}^{\text{salient}}=\{n_{j}\mid\frac{n_{j}}{\max(n_{j})}>0.7\},\forall j\in\text{Layer}$
@@
Input sequence-based: To identify neurons that are salient in the context of the input sequence, we calculate the total attribution over the entire input sequence by summing the attributions across all input tokens. We select the neurons that have an attribution score >0.7>0.7 after normalization. This approach ensures that the selected neurons reflect their contributions to the overall meaning of the input, rather than being limited to the most attributed token only. Given as:
@@
$n_{j}^{\text{salient}}=\{n_{j}\mid\frac{s_{j}}{\max(s_{j})}>0.7\},\forall j\in{Layer}$
@@
$n_{j}^{\text{salient}}=\{n_{j}\mid\frac{m_{j}}{\max(m_{j})}>0.7\},\forall j\in{Layer}$
@@ 2.4 Neuron’s Activations
Since quantization reduces weight precision, it may increase the number of insignificant neurons. To identify them, we follow voita2023neuronslargelanguagemodels, defining dead neurons as those whose activations remain consistently near zero across the dataset.
@@ 2.4.1 Dead/Insignificant Neurons
voita2023neuronslargelanguagemodels observed that the number of dead neurons increases with the growth of a model’s size. Their analysis of the OPT language model family, which uses the ReLU activation function, shows that over 70% of neurons in some layers are dead. We hypothesize that quantization, by reducing the precision of weights, may contribute to an increase in the number of dead neurons in the network.
Apart from ReLU, other activation functions such as GELU DBLP:journals/corr/HendrycksG16 and SiLU DBLP:journals/corr/ElfwingUD17 may not produce activation values that are exactly zero. To generalize the concept of dead neurons for these activation functions, we define a threshold of 0.1-0.1 to $0.1$, categorizing neurons as dead if their activation values consistently remain within this range across the dataset. For different activation functions, we define dead neurons as follows:
@@ 2.5 Correlation Analysis
We hypothesize that a low-precision quantization may cause more neurons to represent identical information, i.e., as precision is reduced, high precision neuron values may map to the same low precision value. Similar to DBLP:journals/corr/abs-2004-04010, we calculate the Pearson correlation of neurons at a layer to identify neurons representing similar information. The Pearson correlation is given by:
@@ 3.1 Datasets
We consider five datasets in this study: BoolQ clark2019boolqexploringsurprisingdifficulty, the Jigsaw Toxicity dataset cjadams_2017, Physical Interaction: Question Answering (PIQA) PiQA, Hellaswag Hellaswag and IMDB sentiment classification sentiment_dataset.
We select a random subset of the each dataset for our experiment. More specifically, we used 10k samples from a combination of train and validation sets of BoolQ, 9k samples from the Toxicity train set, 1,838 validation samples from PIQA, 5,000 validation samples from Hellaswag, and the IMDB training set. Instruction-tuned samples for datasets is available in Appendix A.
@@ 3.2 Models
The primary models analyzed in our study are Phi-2 phi2, Llama-2 7B touvron2023llama2openfoundation, Qwen 2.5 3B and 7B qwen2025qwen25technicalreport, and Mistral-7B jiang2023mistral7b.
@@
Since our subject models employ GELU (Phi-2) and SiLU (Llama, Qwen, Mistral) as activation functions, which do not produce exact zero activations, we include the OPT-6.7B model from the OPT family zhang2022optopenpretrainedtransformer to assess the behavior of ReLU activations for comparison. This model utilizes a decoder-only architecture similar to other subject models.
@@ 3.3 Quantization Configurations
To perform comparative analysis across models under different quantization settings, we employed two widely-used quantization techniques: 4-bit dettmers2023qloraefficientfinetuningquantized and 8-bit dettmers2022llmint88bitmatrixmultiplication. Models are quantized using bitsandbytes config through Huggingface transformers. Table 1 shows the hyperparameters used for quantization.
@@ 3.4 Attribution Technique
We use Integrated Gradients sundararajan2017axiomatic using Captum kokhlikyan2020captum to find salient neurons in a network.
@@ 4.2.2 Calibration Analysis
The seemingly random fluctuations in ACE scores, particularly for certain model, dataset or weight precision combinations, could stem from several underlying factors. Differences in model pretraining objectives or tokenization strategies may contribute to how calibration responds to quantization. Although quantization may introduce fluctuations in ACE, the difference is not substantial to undermine it’s reliability, even often yielding improved calibration relative to full-precision variant.
@@ 4.3 Effect on the Contribution of Neurons to Model Predictions
Overall, the number of salient neurons serves as a proxy for how distributed or localized the decision-making process is within the network durrani2024discoveringsalientneuronsdeep. Full precision models tend to use fewer, neurons when they are smaller. In contrast, in larger models, full precision can enable richer and more distributed neuron contribution.
4.4 Effect on the number of “dead neurons”
As shown in Table 3, quantization causes only a minor change in the count of dead neurons. The trend across quantization seems to be consistent, as the number of dead neurons remains almost similar between quantized and full-precision models.
The pattern of higher neurons in initial layer in Phi-2 and OPT-6.7B likely reflects the role of initial layers in learning sparse, low-level features, while later layers capture higher-level contextual features dalvi2022discoveringlatentconceptslearned; voita2023neuronslargelanguagemodels. We hypothesize that the consistently low count of dead neurons among Llama, Qwen and Mistral is due to the use of the SiLU activation function.
@@ 4.5 Effect on the Redundancy of Neurons
As identified in the works of DBLP:journals/corr/abs-2004-04010 language models can maintain 97% of accuracy while using only 10% of the neurons. This finding is valuable for model pruning. We investigate whether quantization leads to higher redundancy. Due to the substantial computational requirements, our analysis was limited to the Phi-2 and Llama-2-7B models, using activations from BoolQ.
@@ 4.5.1 Correlation Analysis
Figures 4 shows neuron pairs count corresponding to correlation scores for 4-bit, 8-bit and full-precision variants of Phi-2 and Llama-2-7B. The X-axis highlights the different correlation score bins ranging from 0.3-0.4 to 0.9-1.0. This binning process helps to clearly observe the redundant neuron pairs count across all the layers. The Y-axis shows the count of neuron pairs that fall in that bin. Notice that the count is given for neuron pairs across all the layers, as our main focus is to observe the effect on redundancy of neurons within quantizations. For clarity in visualizing highly correlated neuron pairs, we excluded the 0.0–0.1, 0.1–0.2, and 0.2–0.3 correlation bins from the bar graph, since these bins contained similar numbers of uncorrelated neurons across different quantizations.
@@
We hypothesize that the difference in redundancy between Phi-2 and Llama-2-7B is likely due to difference in the number of dead neurons (Table 3). For instance, in the initial layer, Phi-2 has over 20% dead neurons, whereas Llama-2-7B exhibits nearly none. This variation in activations between the two models can likely be attributed to their different activation functions, GELU in Phi-2 and SiLU in Llama-2-7B.
@@ 5 Related Work
Quantization Techniques. Quantization 720541 is used to reduce the memory requirement by reducing the size of weight and/or activation and increasing the inference time of a model jacob2017quantizationtrainingneuralnetworks; DBLP:journals/corr/abs-2103-13630.
Quantization-aware training (QAT) is costly and uses re-training of a model on a dataset to maintain accuracy liu2023llmqatdatafreequantizationaware; du2024bitdistillerunleashingpotentialsub4bit; dettmers2023qloraefficientfinetuningquantized; kim2023memoryefficientfinetuningcompressedlarge.
Post-training quantization quantizes models without any additional finetuning of the model with a limited dataset, but also suffers from performance issues banner2019posttraining4bitquantizationconvolution; DBLP:journals/corr/abs-2001-00281. In case of LLM’s Post Training Quantization can be of 3 types: i) Weight-Only Quantization park2024lutgemmquantizedmatrixmultiplication; frantar2023gptqaccurateposttrainingquantization; chee2024quip2bitquantizationlarge; lin2024awqactivationawareweightquantization, ii) Weight-Activation Quantization NEURIPS2022_adf7fa39; yuan2023rptqreorderbasedposttrainingquantization; 10.1145/3579371.3589038; wei-etal-2023-outlier, and iii) KV Cache Quantization hooper2024kvquant10millioncontext; yue2024wkvquantquantizingweightkeyvalue.
Quantization Analysis and Interpretation. A number of works interpret models in their ability to learn language phenomenon such as morphology belinkov-etal-2017-neural and syntax arps-etal-2022-probing; hall-maudslay-cotterell-2021-syntactic; hupkes_diagnostic. These works often do not relate the representation of linguistic phenomenon to end performance of the model probing_shortcomings. In this work, we aim to experiment with a diverse set of interpretation methods which can be correlated to the performance of the model.
DBLP:journals/corr/abs-2111-08163 explores confidence and calibration relation between quantized and full-precision model by using symmetric quantization. proskurina-etal-2024-quantization shows quantization improves calibration in LLMs using GPTQ. Some literature exlores interpretation withing quantized model for vision model Norrenbrock_Rudolph_Rosenhahn_2024; Xpression; QuantizedAndInterpretable; rezabeyk2024saliencyassistedquantizationneural; 10755934.